# Measurement of the Crystallization and Phase Transition of Niobium Dioxide Thin-Films for Neuromorphic Computing Applications Using a Tube Furnace Optical Transmission System

### v1

@@ II Experimental
Grazing Incidence X-ray Diffraction (GI-XRD) was performed on a Bruker D8 Advance which uses a Cu-K X-ray source. A grazing incidence angle of as well as a slight off-axis rotation of the sample was used to improve the signal to noise ratio of the NbO2 thin films. All samples were measured from to 2 at a rate of 20 s per for a total of 7.64 hr per sample.
@@ IV Results and Discussion
On sapphire, a stronger interaction between the substrate and film has been observed. This is apparent in the orange XRD plot, which is taken at a single, fixed azimuthal angle (ie – sample rotation was not possible in our XRD instrument). We therefore presume that the interaction between the sapphire surface and the NbO2film results in an epitaxial relationship developing during the crystallization process, and that the subsequent NbO2film is textured.

### v2

@@ II Experimental
Grazing incidence X-ray diffraction (GI-XRD) was performed using a Rigaku Smartlab with 9 kW copper rotating anode source and incident and receiving soller slits. The incident angle was fixed at and 2 was scanned from to degrees in steps of .
@@ III Optical Transmission Instrument
Prior to operation of the system, we measured each fixed wavelength for 120 s on 1 s intervals in order to estimate the noise in our optical system. The tube furnace was kept at room temperature during this measurement. The plot of intensity vs. time can be found in the Supplemental section. Across all of the wavelengths, the standard deviation of the optical signal was less than 0.5% of the mean value.
@@ IV Results and Discussion
On sapphire, a stronger interaction between the substrate and film has been observed. This is apparent in the orange XRD plot, for which the relative intensity of the (400) peak compared to the other peaks is greater than that observed for the film on the fused silica substrate, indicating an increase in preferential orientation of the crystalline structure. Note that the absence of the amorphous feature at 2 , which has been previously observed in as-grown Nb2O5(see ref [19]), may be the result of a lower critical angle compared to the film on fused silica (see supplemental figure 3). The lower critical angle suggests that the GIXRD measurement is less surface sensitive, and therefore diminishing the